Two cars depart from the same spot for the same destination. One travels at 60 miles per hour, the other at 65 miles per hour. The faster car arrives an hour before the other car. How far did they travel?
PLEASE HELP!
: Two cars depart from the same spot for the same destination. One travels at 60 miles per hour, the other at 65 miles per hour. The faster car arrives an hour before the other car. How far did they travel?
PLEASE HELP!

Answer by Alan3354(1209)

Two cars depart from the same spot for the same destination. One travels at 60 miles per hour, the other at 65 miles per hour. The faster car arrives an hour before the other car. How far did they travel?
---------------------------------
dist = rate x time
For the faster car,
d = 65t
For the slower car,
d = 60(t+1)
Since the distance is the same,
65t = 60(t+1) (see that?)
65t = 60t + 60
5t = 60
t = 12 hours
--------
Now, d = 65t
d = 780 miles
13*60 also = 780 miles, so it's correct.